Joe has a collection of nickels and dimes that's worth $6.00. If the number of dimes were doubled and the number of nickels were increased by 6, the value of the coins would be $9.90. How many dimes does he have?
Answer by richwmiller(17219)

5n+10d=600
5(n+6)+10*(2d)=990
d=36 n=48
check
5*48+10*36=600
240+360=600
600=600
ok
check
5*(54)+10*72=990
270+720=990
990=990
ok